Calculate the area of a triangle whose vertices are given by (3, −1, 2), (1, −1, 2) 

and (4, −2, 1).


1
Expert's answer
2021-02-04T10:20:14-0500

We must find the area of a triangle whose vertices are given by (3, −1, 2), (1, −1, 2) 

and (4, −2, 1).

A(3, -1, 2); B(1, -1, 2); C(4, -2, 1)

"S = \delta ABC = \large\frac{1}{2}""|\overrightarrow{AB}*\overrightarrow{AC} |"

"\overrightarrow{AB} = -2\overrightarrow{i}" "\overrightarrow{AC} =\overrightarrow{i}-\overrightarrow{j}-\overrightarrow{k}"

"\overrightarrow{AB}*\overrightarrow{AC} =" "\space\space\space \space\space\space \space\space\space \space\space\space \overrightarrow{i} \space\space\space \overrightarrow{j}\space\space\space \overrightarrow{k}"

1 -1 -1 "\to 2\overrightarrow{j}-2\overrightarrow{k}"

-2 0 0

"Area = \frac{1}{2}|2\overrightarrow{j}-2\overrightarrow{k}| = \frac{1}{2}\sqrt{(2)^2+(-2)^2}=\sqrt{2}"

Answer "\sqrt{2}" sq. units